OpenStudy (anonymous):

A car moving with a velocity of 20 meters/second has 1.8 * 105 joules of kinetic energy. What is the mass of the car?

OpenStudy (shane_b):

\[KE = \frac{1}{2}mv^2\]\[1.8x10^5J=\frac{1}{2}m(20m/s)^2\]Now just solve for m...

OpenStudy (shane_b):

If you assistance taking it from there please ask.